# E-THER: A PCT-Grounded Dataset for Benchmarking Empathic AI

### v1

@@ Abstract
A prevalent shortfall among current empathic AI systems is their inability to recognize when verbal expressions may not fully reflect underlying emotional states. This is because the existing datasets, used for the training of these systems, focus on surface-level emotion recognition without addressing the complex verbal-visual incongruence (mismatch) patterns useful for empathic understanding. In this paper, we present E-THER, the first Person-Centered Therapy-grounded multimodal dataset with multidimensional annotations for verbal-visual incongruence detection, enabling training of AI systems that develop genuine rather than performative empathic capabilities. The annotations included in the dataset are drawn from humanistic approach, i.e., identifying verbal-visual emotional misalignment in client-counsellor interactions - forming a framework for training and evaluating AI on empathy tasks. Additional engagement scores provide behavioral annotations for research applications. Notable gains in empathic and therapeutic conversational qualities are observed in state-of-the-art vision-language models (VLMs), such as IDEFICS and VideoLLAVA, using evaluation metrics grounded in empathic and therapeutic principles. Empirical findings indicate that our incongruence-trained models outperform general-purpose models in critical traits, such as sustaining therapeutic engagement, minimizing artificial or exaggerated linguistic patterns, and maintaining fidelity to PCT theoretical framework.
@@ I Introduction
Empathic dialogue generation is a central challenge for human–AI interaction. While large language models (LLMs) and vision–language models (VLMs) have advanced open-domain conversation, systems still tend to rely on surface regularities in text, producing responses that appear empathic without deeper situational grounding [1, 2]. In applied communication settings, this gap is often described as performative empathy—language that signals care but does not reflect nuanced understanding [3].
A key source of nuance arises from multimodal inconsistency: what people say can diverge from how they present nonverbally. Counseling and communication theories describe such verbal–visual incongruence as diagnostically meaningful [4, 5]. Existing empathy and emotion resources—spanning text-only datasets [6, 7] and multimodal corpora [8, 9, 10, 11]—provide valuable foundations for response generation and emotion recognition. However, systematic annotation and modeling of verbal–visual incongruence for empathic response generation remain comparatively underexplored. In parallel, engagement level (e.g., low vs. high client engagement) is known to shape effective conversational strategies [12], yet many pipelines treat empathic behaviors as context-invariant despite evidence that engagement awareness can enhance interaction quality [13].

### v2

# E-THER: A Multimodal Dataset for Empathic AI - Towards Emotional Mismatch Awareness

@@ Abstract
A prevalent shortfall among current empathic AI systems is their inability to recognize when verbal expressions may not fully reflect underlying emotional states. This is because the existing datasets, used for the training of these systems, focus on surface-level emotion recognition without addressing the complex verbal-visual incongruence (mismatch) patterns useful for empathic understanding. In this paper, we present E-THER, the first Person-Centered Therapy (PCT)-grounded multimodal dataset with multidimensional annotations for verbal-visual incongruence detection, enabling training of AI systems to advance towards realistic rather than performative empathic capabilities. The annotations included in the dataset are drawn from humanistic approach, i.e., identifying verbal-visual emotional misalignment in client-counsellor interactions - forming a framework for training and evaluating AI on empathy tasks. Additional engagement scores provide behavioral annotations for research applications. Notable gains in empathic and therapeutic conversational qualities are observed in state-of-the-art vision-language models, such as IDEFICS and VideoLLAVA, using evaluation metrics following empathic and therapeutic principles. Empirical findings indicate that our incongruence-trained models outperform general state-of-the-art models in critical traits, such as sustaining therapeutic engagement, minimizing artificial or exaggerated linguistic patterns, and maintaining fidelity to PCT theoretical framework.
@@ I Introduction
Empathic dialogue generation is a central challenge for human–AI interaction. While large language models (LLMs) and vision–language models (VLMs) have advanced open-domain conversation, systems still tend to rely on surface regularities in text, producing responses that appear empathic without deeper situational grounding [1, 2]. In applied communication settings, this gap is often described as performative empathy - language that signals care but does not reflect nuanced understanding [3].
A key source of nuance arises from multimodal inconsistency: what people say can diverge from how they present nonverbally. Counseling and communication theories describe such verbal–visual incongruence as diagnostically meaningful [4, 5]. Existing empathy and emotion resources, spanning text-only datasets [6, 7] and multimodal corpora [8, 9, 10, 11], provide valuable foundations for response generation and emotion recognition. However, systematic annotation and modeling of verbal–visual incongruence for empathic response generation remain comparatively underexplored. In parallel, engagement level (e.g., low vs. high client engagement) is known to shape effective conversational strategies [12], yet many pipelines treat empathic behaviors as context-invariant despite evidence that engagement awareness can enhance interaction quality [13].